On December 27, 2018, Sunland Windows purchased a piece of equipment for $109,000. The estimated useful life of the equipment is either three years or 50,000 units, with a residual value of $7,000. The company has a December 31 fiscal year end and normally uses straight-line depreciation. Management is considering the merits of using the units-of-production or diminishing-balance method of depreciation instead of the straight-line method. The actual numbers of units produced by the equipment were 12,000 in 2019, 19,500 in 2020, and 17,500 in 2021. The equipment was sold on January 5, 2022, for $15,000.

Calculate the depreciation for the equipment for 2019 to 2021 under

1. the straight-line method
2. the diminishing-balance method, using a 40% rate; and
3. units-of-production


(Round depreciable amount per unit to 3 decimal places, e.g. 1.252 and the final answers to 0 decimal places, e.g. 126.)

Depreciation Expense
Straight-Line Diminishing-Balance Units-of-Production
2019 $ $ $
2020
2021

eTextbook and Media

  

  

Calculate the gain or loss on the sale of the equipment under each of the three methods. (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Gain (loss) on sale
1. Straight-line $                                                                       gainloss
2. Diminishing-balance $                                                                       gainloss
3. Unit-of-production $                                                                       gainloss

eTextbook and Media

  

  

Calculate the total depreciation expense plus the loss on sale (or minus the gain on sale) under each of the three depreciation methods.

Net Expense
1. Straight-line $
2. Diminishing-balance $
3. Unit-of-production $
Answer #1

Under Straight line method

Depreciation expense per year= (Original cost-Residual value)/Estimated useful life

= ($109000-7000)/3

= $34000

Year Depreciation expense Accumulated depreciation Book value at the end of the year
2019 $34000 $34000 (109000-34000)= 75000
2020 34000 (34000+34000)= 68000 (75000-34000)= 41000
2021 34000 (68000+34000)= 102000 (41000-34000)= 7000

Under Diminishing-balance method

Depreciation rate= 40%

Year Depreciation expense Accumulated depreciation Book value at the end of the year
2019 (109000*40%)= $43600 $43600 (109000-43600)= 65400
2020 (65400*40%)= 26160 (43600+26160)= 69760 (65400-26160)= 39240
2021 (39240*40%)= 15696 (69760+15696)= 85456 (39240-15696)= 23544

Under Units-of-production method

Depreciation expense per year= (Original cost-Residual value)/Estimated units

= ($109000-7000)/50000

= $2.04 per year

Year Depreciation expense Accumulated depreciation Book value at the end of the year
2019 (12000*$2.04)= 24480 $24480 (109000-24480)= 84520
2020 (19500*$2.04)= 39780 (24480+39780)= 64260 (84520-39780)= 44740
2021 (17500*$2.04)= 35700 (64260+35700)= 99960 (44740-35700)= 9040
Depreciation Expense
Straight-Line Diminishing-Balance Units-of-Production
2019 $34000 $43600 $24480
2020 34000 26160 39780
2021 34000 15696 35700

Calculate the gain or loss on the sale of the equipment under each of the three methods

Straight-Line Diminishing-Balance Units-of-Production
Book value at the end of 2019 $7000 $23544 $9040
Sales prices 15000 15000 15000
Loss (gain) on sale of equipment $(8000) $8544 $(5960)
Gain Loss Gain

Calculate the total depreciation expense plus the loss on sale (or minus the gain on sale) under each of the three depreciation methods.

Straight-Line Diminishing-Balance Units-of-Production
Total depreciation expense $102000 $85456 $99960
Loss (gain) on sale of equipment (8000) 8544 (5960)
Net expense $94000 $94000 $94000